PREX Meeting Notes, Mar 11, 2010

Present: Bob Michaels, Ole Hansen, Mindy Jen, Luis Mercado,
         Krishna Kumar, Megan Friend, Gregg Franklin, Brian Quinn, 
         Richard Holmes, Sasha Glamazdin, Paul King,
         Alexandre Camsonne, Zafar Ahmed, Diancheng Wang,
         Jon Wexler, Seamus Riordan, Guido Urciuoli,
         Doug Higinbotham, Abdurahim Rakham, Al Tobias,
         Dustin McNulty, Tim Holmstrom, Kiadtisak Saenboonruang

Shifts start Wed Mar 17 according to present schedule.

Next week start daily meetings in 2nd floor counting room 
on Tuesday Mar 16.  Each day at 3 pm.

Luis / DAQ

    DAQ is basically ready but will swap out 6 ADC16 for ADC18 in CH.
    Updated the ADC list
    Questions re: panFFB speed.  Richard Holmes to investigate

Compton -- Abdurahim and Al 

    Optics ready.  Beam alignment done.
    All hardware in place.  Polarization measurements done.
    Got better polarization at CIP after replacing mirrors.
    Polarization transfer function: measurements done but
    "might be incomplete" (this needs to be clarified).
    New cavity mirrors: gain = 4500
    Pumping out vacuum now, will bake out soon.
    Tried to align light in cavity, tried to lock -- didn't work.
    Problems with locking listed -- not sure I (the author of
    these notes) understood it, so I won't write it here.
    Still to do: calibrations of power meter and other stuff,
    some EPICS control software.
Megan / Brian Q. -- Compton DAQ

    It's ready !



     L-bracket platform on both HRS
     X-Y sliders hooked to portserver
     Supports for scint.
     Cabling for quartz
     Hardware for detector mounting
     Very active installation schedule this weekend


     Cables ready
     Need to setup trigger system (function generator)
     Need BMW mods

Ole -- Gems

     Chambers from INFN and UVa
     5 chambers on site, 6th arrives tomorrow from UVa
     All being installed in 1 HRS
     KK asks: Can we point from VDCs to test Gems ? (ans: Yes)

Sasha -- Moller

     Vacuum problem fixed 
     Will cool magnet
     Meetings w/ MCC re: Moller procedures and apparatus.
     FSD requirements ironed out.
     Meeting re: pulsed mode (expert mode).
        Will be "slightly invasive" to hall B (I'm still
             not sure what this means)

     SU scint:  noisy at 900V, ok at 700V.
     testing with cosmics
     Dustin asks: using a x10 amplifier ?  (ans: Yes)
     Checking connections for cavities and Lumis
     KK comment: will review all noise issues and see what
       needs to be done.  Need hel-corr x-talk studies, etc.


     Working with Megan, learning about Compton


     RCs will need to undergo training to be able to take
     the load off techs at nights or odd times.  The protocol
     for calling techs will be clarified; essentially all
     calls at odd hours for routine fixes like resetting Q1 must go 
     through the RC if the shift workers don't know what to do.

 Next meeting   Tues Mar 16 at 3 pm   2nd floor CH

Index of Meetings